OpenStudy (anonymous):

Help with Algebra 2?

OpenStudy (muscrat123):

maybe :)

OpenStudy (anonymous):

haha :) Hold on please

OpenStudy (muscrat123):

put the question in the question section though plz - edit the ?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Jean and Mark are going to fill a pool with 2 different sized hoses. Jean can fill the pool in 8 hours, while Mark can complete it in 12 hours. Their supervisor thinks that the job will take 10 hours to complete if they work together. Explain each step in solving this equation, and determine if the supervisor is correct or not.

OpenStudy (anonymous):

The "edit question" isn't working. I typed my question but it won't let me submit it.

OpenStudy (muscrat123):

oh ok.

OpenStudy (muscrat123):

so jean fills 1/8 of the pool per hour. do u understand that?

OpenStudy (muscrat123):

@KaseyClinta

OpenStudy (anonymous):

yes!

OpenStudy (muscrat123):

ok. so then for mark.

OpenStudy (anonymous):

Alright:)

OpenStudy (welshfella):

Find rate of filling for each person Jean fills pool on 8 hrs IN hour she fills 1/8 of the pool Mark fills 1/12 in 1 hour together let them they take x hours so we have 1/8 + 1/12 = 1/x Solve for x

OpenStudy (muscrat123):

does that help? ^^^^^^^^^^^^^^^^^^^^^^^

OpenStudy (anonymous):

yes. So the LCM would be 24x?

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

So I multiply each fraction by that correct?

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

alright, hold on

OpenStudy (muscrat123):

youre doing great!

OpenStudy (anonymous):

3x/24x + 2x/24x =24/24x

OpenStudy (anonymous):

Thanks:)

OpenStudy (anonymous):

@muscrat123

OpenStudy (muscrat123):

yes?

OpenStudy (anonymous):

what do I do now?

OpenStudy (anonymous):

@Kash_TheSmartGuy @Jhannybean @freckles

OpenStudy (anonymous):

5x=24 ??

OpenStudy (welshfella):

well each fraction has a denominator of 24x yes thats right

OpenStudy (anonymous):

So what would be the answer?

OpenStudy (welshfella):

divide both sides by 5 to get x = ?

OpenStudy (anonymous):

0.208

OpenStudy (welshfella):

24 / 5 = 0.208????

OpenStudy (anonymous):

4.8 @welshfella

OpenStudy (welshfella):

yea that's it 4 .8 hours

OpenStudy (anonymous):

So then the guy would be wrong, right?

OpenStudy (welshfella):

right

OpenStudy (anonymous):

:)